# Finite element approximation of the stationary Navier-Stokes problem with non-smooth data

### v1

@@ 3 Error analysis of approximating boundary data with regular functions
In this section, we obtain estimates of the error involved in approximating our problem (2) with one that uses regular boundary data. To achieve this, we show some useful descomposition of the solution of (2), into a regular part and a non-regular part.
@@
There are some proposals for $\mathbf{g}_{\varepsilon}$ that can be found in [1, 4, 3] which we will analyze in the numerical experiments section.
In [9] the authors present the following decomposition of the very weak solution of (2), which will be very useful for our objectives.
@@
Let $\Omega\subset\mathbb{R}^{2}$ be a domain that satifies Assumption 1. Given $\varepsilon>0$, $\mathbf{f}\in\mathbf{H}^{-1}(\Omega)$ and $\mathbf{g}_{\varepsilon}\in\mathbf{H}^{1/2}(\Gamma)$ a regularization of $\mathbf{g}$, there exists u very weak solution of (2) which can be decomposed as $\textit{{u}}=\textit{{u}}_{\varepsilon}+\textit{{v}}_{\varepsilon}$, where $\textit{{u}}_{\varepsilon}\in\mathbf{L}_{\Gamma}$ is a very week solution of
@@
$-\nu\int_{\Omega}(\textit{{u}}_{\varepsilon}+\textit{{v}}_{\varepsilon})\cdot\Delta\phi-\int_{\Omega}((\textit{{u}}_{\varepsilon}+\textit{{v}}_{\varepsilon})\cdot\nabla)\phi\cdot(\textit{{u}}_{\varepsilon}+\textit{{v}}_{\varepsilon})$ $=\int_{\Omega}\mathbf{f}\cdot\phi-\int_{\Gamma}\mathbf{g}\cdot\frac{\partial\phi}{\partial\mathbf{n}}\quad\forall\phi\in\mathbf{X}$
@@
Let $\Omega\subset\mathbb{R}^{2}$ be a domain that satifies Assumption 1. Given $\varepsilon>0$ and let $(\textit{{v}},q)\in\mathbf{H}^{1}(\Omega)\times L^{2}_{0}(\Omega)$ be the weak solutions of (3). If
@@
with $\hat{\mathbf{f}}=\mathbf{f}+\nu\Delta\Psi-(\Psi\cdot\nabla)\Psi$, where $\Psi\in\mathbf{H}^{1}(\Omega)$ is the regular extension function of $\mathbf{g}_{\varepsilon}$ provided by lemma 2.2.
Then, there exists a positive constant $C$ such that
$|\textit{{v}}_{\varepsilon}-\textit{{v}}|_{H^{1}(\Omega)}\leq C\frac{C_{1}\|\mathbf{g}-\mathbf{g}_{\varepsilon}\|_{L^{2}(\Gamma)}}{1-C_{1}(1/\nu)\|\mathbf{g}-\mathbf{g}_{\varepsilon}\|_{L^{2}(\Gamma)}}$
@@
For $\mathbf{g}_{\varepsilon}\in\mathbf{H}^{1/2}(\Gamma)$, by Lemma 2.2 we know that there exists a function $\Psi\in\mathbf{H}^{1}(\Omega)$ such that $\textup{div}\ \Psi=0$ and $\Psi|_{\Gamma}=\mathbf{g}_{\varepsilon}$. Following the ideas of the proof of the Theorem 2.3, we can write $\textit{{v}}_{\varepsilon}=\hat{\textit{{v}}}_{\varepsilon}+\Psi$ and $\textit{{v}}=\hat{\textit{{v}}}+\Psi$ with $\hat{\textit{{v}}}_{\varepsilon},\hat{\textit{{v}}}\in\mathbf{H}^{1}_{0}(\Omega)$.
Given $\phi\in\mathbf{H}^{1}_{0}(\Omega)$, with $\textup{div}\ \phi=0$, a test function for the variational problem of (3) and (10),
@@
$\left\{\begin{array}[]{rlc}a(\textit{{v}},\textit{{w}})+c(\textit{{v}},\textit{{v}},\textit{{w}})-b(\textit{{w}},q)&=\int_{\Omega}\mathbf{f}\cdot\textit{{w}}&\forall\textit{{w}}\in\mathbf{H}^{1}_{0}(\Omega)\\ \textup{div}\ \textit{{v}}&=0&\textup{in }\,\Omega\\ \textit{{v}}&=\mathbf{g}_{\epsilon}&\textup{on }\,\Gamma\end{array}\right$. (14)
@@
$\nu|\hat{\textit{{v}}}_{\varepsilon}-\hat{\textit{{v}}}|^{2}_{H^{1}(\Omega)}$ $\leq\mathcal{C}_{NL}|\hat{\textit{{v}}}_{\varepsilon}-\hat{\textit{{v}}}|^{2}_{H^{1}(\Omega)}|\hat{\textit{{v}}}|_{H^{1}(\Omega)}+\varepsilon|\hat{\textit{{v}}}_{\varepsilon}-\hat{\textit{{v}}}|^{2}_{H^{1}(\Omega)}+C\|\textit{{u}}_{\varepsilon}\|_{L^{4}(\Omega)}\|\textit{{v}}_{\varepsilon}\|_{H^{1}(\Omega)}\|\hat{\textit{{v}}}_{\varepsilon}-\hat{\textit{{v}}}\|_{H^{1}(\Omega)}$
$\leq|\hat{\textit{{v}}}_{\varepsilon}-\hat{\textit{{v}}}|^{2}_{H^{1}(\Omega)}\frac{\mathcal{C}_{NL}\|\hat{\mathbf{f}}\|_{-1}}{\nu-\varepsilon}+\varepsilon|\hat{\textit{{v}}}_{\varepsilon}-\hat{\textit{{v}}}|^{2}_{H^{1}(\Omega)}+C\|\textit{{u}}_{\varepsilon}\|_{L^{4}(\Omega)}\|\textit{{v}}_{\varepsilon}\|_{H^{1}(\Omega)}|\hat{\textit{{v}}}_{\varepsilon}-\hat{\textit{{v}}}|_{H^{1}(\Omega)}$
@@
Let $\Omega\subset\mathbb{R}^{2}$ be a domain that satifies Assumption 1. Given $\varepsilon>0$, $\mathbf{g}\in\mathbf{L}^{2}(\Gamma)$ and $\mathbf{g}_{\varepsilon}\in\mathbf{H}^{1/2}(\Gamma)$ such that $\|\mathbf{g}-\mathbf{g}_{\varepsilon}\|_{L^{2}(\Gamma)}<\varepsilon$. If $\textit{{u}}\in\mathbf{L}_{\Gamma}$ is very weak solution of (2) and v is a weak solution of (3) then
@@
Since u is a very weak solution, by the descomposition given in theorem 3.1 we can write $\textit{{u}}=\textit{{u}}_{\varepsilon}+\textit{{v}}_{\varepsilon}$. Hence, by (8) and the estimates (11) and (15) we get
@@ 4 Finite element approximation and a priori estimates
Let $\Omega$ be a domain satisfying Assumption 1 and with a $C^{k+1}$ piecewise smooth boundary, with $k\geq 1$ sufficiently large to fulfill our requirements. Let $\{\mathcal{T}_{h}\}$ be a partition of $\Omega$ which represent convincingly the domain, i.e., we use curved triangles along the curved part of the boundary (as, for example, the curved triangles introduced by Zlamal in [15]) and classical triangles in the rest of the domain. We use the same notation to call the curved elements as for classical triangles.
Let ${\mathcal{T}}_{h}$ be a family of triangulations of $\Omega$ such that any two triangles in ${\mathcal{T}}_{h}$ share at most a vertex or an edge and each element $T\in{\mathcal{T}}_{h}$. We denote by $h_{T}$ and $\theta_{T}$ the greatest side and smallest angle of the triangle $T$. We assume that the family of triangulations $\{{\mathcal{T}}_{h}\}$ satisfies a minimum angle condition, i.e., there exists a constant $\theta_{0}>0$ such that $\theta_{T}\geq\theta_{0}$, for any $T\in{\mathcal{T}}_{h}$. We also denote by $h=\max_{T\in{\mathcal{T}}_{h}}h_{T}$, i.e., the maximum diameter of the elements in $\mathcal{T}_{h}$.
@@
    For each $\textit{{w}}\in\mathbf{H}^{2}(\Omega)\cap\{\phi\in\mathbf{H}^{1}_{0}(\Omega):\textup{div}\ \phi=0\}$ and $q\in H^{1}(\Omega)\cap L^{2}_{0}(\Omega)$, there exist approximations $\pi_{h}(\textit{{w}})\in\mathbf{V}_{h}$ and $\rho_{h}(q)\in Q_{h}$, such that
    $|\textit{{w}}-\pi_{h}(\textit{{w}})|_{H^{1}(\Omega)}\leq Ch^{r}\|\textit{{w}}\|_{H^{2}(\Omega)},\quad\|q-\rho_{h}(q)\|_{L^{2}(\Omega)}\leq Ch^{r}\|q\|_{H^{1}(\Omega)},$
@@
    The following inverse inequality
    $|\textit{{w}}_{h}|_{H^{1}(\Omega)}\leq Ch^{-1}\|\textit{{w}}_{h}\|_{L^{2}(\Omega)},\quad\textit{{w}}_{h}\in\mathbf{V}_{h}.$
  • A3:
@@
There are several families of elements present in the literature that satisfy conditions A1, A2, and A3 (see, for example, [16, 17, 18, 19, 20, 21] and the references therein). From now on, let us consider $\mathbf{g}_{\varepsilon}=\mathbf{g}_{h}$, i.e. a regular approximation of $\mathbf{g}$, such that $\mathbf{g}_{h}$ is the trace of a function in $\mathbf{V}_{h}$ and
@@
$\left\{\begin{array}[]{rll}a(\textit{{u}}_{h},\phi_{h})+c(\textit{{u}}_{h}.\textit{{u}}_{h},\phi_{h})-b(\phi_{h},p_{h})&=\int_{\Omega}\mathbf{f}\cdot\phi_{h}\quad&\forall\phi_{h}\in\mathbf{V}_{h}\cap\mathbf{H}^{1}_{0}(\Omega)\\ b(\textit{{u}}_{h},q_{h})&=0\quad&\forall q_{h}\in Q_{h}\end{array}\right$. (17)
The existence and uniqueness of problem (17) is studied, for example, in [8, 22], for conforming finite element methods that satisfy A1,A2 and A3. In the next section, we consider some possible choices for the discrete spaces $\mathbf{V}_{h}$, $Q_{h}$ and for the function $\mathbf{g}_{h}$.
The following result provides an useful a priori error estimate for the finite element approximation of Navier-Stokes problem with regular data.
@@
Let $\mathbf{f}\in\mathbf{H}^{-1}(\Omega)$, if A1, A2 and A3 hold, we get the following error estimates
@@
Let $\Omega\subset\mathbb{R}^{2}$ be a domain that satisfies Assumption 1 and let $\mathbf{f}\in\mathbf{H}^{-1}(\Omega)$ be and $\mathbf{g}_{h}$ satisfying (16). Let $\textit{{u}}\in\mathbf{L}_{\Gamma}$ be the solution of the problem (2) and $\textit{{u}}_{h}$ be the solution of (17) . If A1, A2 and A3 hold. Then, there exists a constant $C$ such that
@@ 5 Numerical experiments
In this section we discuss some numerical aspects of our approach and present numerical experiments for the classic lid-driven cavity flow problem, which is a typical benchmark problem in computational fluid dynamics [3, 4, 17].
The first issue is to choose finite elements spaces $\mathbf{V}_{h}$ and $Q_{h}$ satisfying conditions A1, A2 and A3. We will consider the following:
  • 1.
@@
  • 3.
    $\mathbf{P}_{1}P_{1}$ stabilized: the method consists of find $(\textit{{u}}_{h},p_{h})\in\mathbf{V}_{h}\times Q_{h}$ with $\textit{{u}}_{h}=\mathbf{g}_{\varepsilon}$ in $\Gamma$, $\mathbf{V}_{h}=(P_{1}(\mathcal{T}_{h}))^{2}$ and $Q_{h}=P_{1}(\mathcal{T}_{h})\cap L^{2}_{0}(\Omega)$ such that
    $\left\{\begin{array}[]{rll}a(\textit{{u}}_{h},\phi_{h})+c(\textit{{u}}_{h}.\textit{{u}}_{h},\phi_{h})-b(\phi_{h},p_{h})&=\int_{\Omega}\mathbf{f}\cdot\phi_{h}&\forall\phi_{h}\in\mathbf{V}_{h}\cap\mathbf{H}^{1}_{0}(\Omega)\\ b(\textit{{u}}_{h},q_{h})-G(p_{h},q_{h})&=0&\forall q_{h}\in Q_{h}\end{array}\right$.
@@
    with $\Pi q|_{T}=\frac{1}{|T|}\int_{T}qdx,\forall T\in\mathcal{T}_{h}$. This method was proposed in [19] for the Stokes problem and studied for the Navier-Stokes problem in [23, 24].
After one of those methods is chosen, another question is how to select an appropriate function $\mathbf{g}_{h}$ that satisfies (16) and the compatibility condition (1). One possibility is to take $\mathbf{g}_{h}$ as the Cartensen interpolator of the function $\mathbf{g}$ [25]. Another option is to take $\mathbf{g}_{h}$ as the Lagrange interpolator but, unfortunately, this choice doesn’t verify (1). In [3, lemma 3.1] the authors propose a modification of the Lagrange interpolator so that (1) holds.
Let $\Omega=[-1,1]\times[-1,1]$, $\mathbf{f}=(0,0)$ and
@@
We will obtain the numerical approximations applying the three finite elements methods described above and choosing $\mathbf{g}_{h}$ as the Lagrange interpolator. We observe that, in this case, the compatibility condition is automatically verifed and furthermore, following [3, Proposition 3.1], we have that
@@
Other numerical aspect to consider is the iterative methods to approximate the discrete solution. The typical algorithm to solve the Navier-Stokes problem is based on a fixed-point iterative algorithm. One possibility, that we will apply in our numerical experiments, is to use a Newton scheme, which has been studied for both the homogeneous [10, 26] and non-homogeneous cases [8, 22].
@@
$a(\textit{{u}}^{n}_{h},\phi_{h})+c(\textit{{u}}^{n}_{h}.\textit{{u}}^{n-1}_{h},\phi_{h})$ $+c(\textit{{u}}^{n-1}_{h}.\textit{{u}}^{n}_{h},\phi_{h})-b(\phi_{h},p^{n}_{h})$
@@
In [8, Proposition 3.4 and Proposition 3.5] it is shown that if $\textit{{u}}^{0}_{h}$ is is close enough to $\textit{{u}}_{h}$, then the sequence defined by (5) is unique and converges quadratically to $\textit{{u}}_{h}$. In other words, there exists a $\delta>0$ such that if $\|\textit{{u}}_{0}-\textit{{u}}_{h}\|_{H^{1}(\Omega)}\leq\delta$, then for $n\geq 1$,
@@
The implementation is based on the FreeFem++ code [27]. For the application of Newton’s method it is necessary to give a stopping criterion given a tolerance value. In our case we use the relative error between two consecutive iterations
$\frac{\|\textit{{u}}^{n}_{h}-\textit{{u}}^{n-1}_{h}\|_{H^{1}(\Omega)}}{\|\textit{{u}}^{n}_{h}\|_{H^{1}(\Omega)}}<tol$
with $tol=10^{-9}$.
@@
Tables 1, 2 and 3 show the graphics of $\log(h)$ vs $\log(e_{L^{4}})$ for $P_{2}P_{1}$, $P^{b}_{1}P_{1}$ and stabilized $P_{1}P_{1}$ elements respectively, achieved with uniform mesh refinement, and also exhibit the experimental order of convergence (eoc) with respect to $h$, which is computed by comparing the errors for two consecutive meshes. We can observe that, in all the cases, the numerical results confirm the theorical prediction.
Table 1: $P_{2}P_{1}$ with uniform mesh refinement.
nv $h$ $e_{L^{4}}$ eoc
289 1/8 0.0952183 -
1089 1/16 0.0673277 0.50004
4225 1/32 0.0476136 0.49993
16641 1/64 0.0336701 0.49991
Table 2: $P^{b}_{1}P_{1}$ with uniform mesh refinement.
nv $h$ $e_{L^{4}}$ eoc
289 1/8 0.173603 -
1089 1/16 0.122749 0.50008
4225 1/32 0.0867935 0.50007
16641 1/64 0.061369 0.50007
Table 3: $P_{1}P_{1}$ stabilization with uniform mesh refinement.
nv $h$ $e_{L^{4}}$ eoc
289 1/8 0.171512 -
1089 1/16 0.121243 0.50041
4225 1/32 0.0857058 0.50042
16641 1/64 0.0605894 0.50039

### v2

@@ 3 Error analysis of approximating boundary data with regular functions
In this section, we obtain estimates of the error involved in approximating our problem (2) with one that uses regular boundary data. To achieve this, we show some useful decomposition of the solution of (2), into a regular part and a non-regular part.
@@
There are some proposals for $\mathbf{g}_{\varepsilon}$ that can be found in [apel2016, duran2020, hamouda2017] which we will analyze in the numerical experiments section.
In [marusic2000] the author present the following decomposition of the very weak solution of (2), which will be very useful for our objectives.
@@
Given $\varepsilon>0$, $\mathbf{f}\in\mathbf{H}^{-1}(\Omega)$ and $\mathbf{g}_{\varepsilon}\in\mathbf{H}^{1/2}(\Gamma)$ a regularization of $\mathbf{g}$, there exists u very weak solution of (2) which can be decomposed as $\textit{{u}}=\textit{{u}}_{\varepsilon}+\textit{{v}}_{\varepsilon}$, where $\textit{{u}}_{\varepsilon}\in\mathbf{L}_{\Gamma}$ is a very week solution of
@@
$-\nu\int_{\Omega}(\textit{{u}}_{\varepsilon}+\textit{{v}}_{\varepsilon})\cdot\Delta\phi-\int_{\Omega}((\textit{{u}}_{\varepsilon}+\textit{{v}}_{\varepsilon})\cdot\nabla)\phi\cdot(\textit{{u}}_{\varepsilon}+\textit{{v}}_{\varepsilon})$ $=\int_{\Omega}\mathbf{f}\cdot\phi$
$-\int_{\Gamma}\mathbf{g}\cdot\frac{\partial\phi}{\partial\mathbf{n}}\quad\forall\phi\in\mathbf{X}$
@@
Given $\varepsilon>0$ and let $(\textit{{v}},q)\in\mathbf{H}^{1}(\Omega)\times L^{2}_{0}(\Omega)$ be the weak solutions of (3). If
@@
with $\hat{\mathbf{f}}=\mathbf{f}+\nu\Delta\Psi-(\Psi\cdot\nabla)\Psi$, where $\Psi\in\mathbf{H}^{1}(\Omega)$ is the regular extension function of $\mathbf{g}_{\varepsilon}$ provided by lemma 2.2. Then, there exists a positive constant $C$ such that
$|\textit{{v}}_{\varepsilon}-\textit{{v}}|_{H^{1}(\Omega)}\leq C\frac{C_{1}\|\mathbf{g}-\mathbf{g}_{\varepsilon}\|_{L^{2}(\Gamma)}}{1-C_{1}(1/\nu)\|\mathbf{g}-\mathbf{g}_{\varepsilon}\|_{L^{2}(\Gamma)}}.$
@@
For $\mathbf{g}_{\varepsilon}\in\mathbf{H}^{1/2}(\Gamma)$, by Lemma 2.2 we know that there exists a function $\Psi\in\mathbf{H}^{1}(\Omega)$ such that $\textup{div}\ \Psi=0$ and $\Psi|_{\Gamma}=\mathbf{g}_{\varepsilon}$. Following the ideas of the proof of the Theorem 2.3, we can write $\textit{{v}}_{\varepsilon}=\hat{\textit{{v}}}_{\varepsilon}+\Psi$ and $\textit{{v}}=\hat{\textit{{v}}}+\Psi$ with $\hat{\textit{{v}}}_{\varepsilon},\hat{\textit{{v}}}\in\mathbf{H}^{1}_{0}(\Omega)$. Given $\phi\in\mathbf{H}^{1}_{0}(\Omega)$, with $\textup{div}\ \phi=0$, a test function for the variational problem of (3) and (10),
@@
$a(\textit{{v}},\textit{{w}})+c(\textit{{v}},\textit{{v}},\textit{{w}})=\int_{\Omega}\mathbf{f}\cdot\textit{{w}}\quad\forall\textit{{w}}\in\mathbf{V}$ (14)
@@
$\nu|\hat{\textit{{v}}}_{\varepsilon}-\hat{\textit{{v}}}|^{2}_{H^{1}(\Omega)}$ $\leq\mathcal{C}_{NL}|\hat{\textit{{v}}}_{\varepsilon}-\hat{\textit{{v}}}|^{2}_{H^{1}(\Omega)}|\hat{\textit{{v}}}|_{H^{1}(\Omega)}+\varepsilon|\hat{\textit{{v}}}_{\varepsilon}-\hat{\textit{{v}}}|^{2}_{H^{1}(\Omega)}$
$+C\|\textit{{u}}_{\varepsilon}\|_{L^{4}(\Omega)}\|\textit{{v}}_{\varepsilon}\|_{H^{1}(\Omega)}\|\hat{\textit{{v}}}_{\varepsilon}-\hat{\textit{{v}}}\|_{H^{1}(\Omega)}$
$\leq|\hat{\textit{{v}}}_{\varepsilon}-\hat{\textit{{v}}}|^{2}_{H^{1}(\Omega)}\frac{\mathcal{C}_{NL}\|\hat{\mathbf{f}}\|_{-1}}{\nu-\varepsilon}+\varepsilon|\hat{\textit{{v}}}_{\varepsilon}-\hat{\textit{{v}}}|^{2}_{H^{1}(\Omega)}$
$+C\|\textit{{u}}_{\varepsilon}\|_{L^{4}(\Omega)}\|\textit{{v}}_{\varepsilon}\|_{H^{1}(\Omega)}|\hat{\textit{{v}}}_{\varepsilon}-\hat{\textit{{v}}}|_{H^{1}(\Omega)}$
@@
Given $\varepsilon>0$ and $\mathbf{g}_{\varepsilon}\in\mathbf{H}^{1/2}(\Gamma)$ such that $\|\mathbf{g}-\mathbf{g}_{\varepsilon}\|_{L^{2}(\Gamma)}<\varepsilon$. Assuming that $\frac{\mathcal{C}_{NL}\|\hat{\mathbf{f}}\|_{-1}}{(\nu-\varepsilon)^{2}}<1$ and $\|\mathbf{g}\|_{L^{2}(\Gamma)}$ is small enough, if $\textit{{u}}\in\mathbf{L}_{\Gamma}$ is very weak solution of (2) and v is a weak solution of (3) then
@@
Since u is a very weak solution, by the decomposition given in theorem 3.1 we can write $\textit{{u}}=\textit{{u}}_{\varepsilon}+\textit{{v}}_{\varepsilon}$. Hence, by (8) and the estimates (11) and (15) we get
@@ 4 Finite element approximation and a priori estimates
Let $\Omega$ be a domain satisfying Assumption 1 and with a $C^{k+1}$ piecewise smooth boundary, with $k\geq 1$ sufficiently large to fulfill our requirements. Let $\mathcal{T}_{h}$ be a partition of $\Omega$ which represent convincingly the domain, i.e., we use curved triangles along the curved part of the boundary (as, for example, the curved triangles introduced by Zlamal in [zlamal1973]) and classical triangles in the rest of the domain. We use the same notation to call the curved elements as for classical triangles.
Let $\{\mathcal{T}_{h}\}$ be a family of triangulations of $\Omega$ such that any two triangles in ${\mathcal{T}}_{h}$ share at most a vertex or an edge and each element $T\in{\mathcal{T}}_{h}$. We denote by $h_{T}$ and $\theta_{T}$ the greatest side and smallest angle of the triangle $T$. We assume that the family of triangulations $\{{\mathcal{T}}_{h}\}$ satisfies a minimum angle condition, i.e., there exists a constant $\theta_{0}>0$ such that $\theta_{T}\geq\theta_{0}$, for any $T\in{\mathcal{T}}_{h}$. We also denote by $h=\max_{T\in{\mathcal{T}}_{h}}h_{T}$, i.e., the maximum diameter of the elements in $\mathcal{T}_{h}$.
@@
    For each $\textit{{w}}\in\mathbf{H}^{2}(\Omega)\cap\{\phi\in\mathbf{H}^{1}_{0}(\Omega):\textup{div}\ \phi=0\}$ and $q\in H^{1}(\Omega)\cap L^{2}_{0}(\Omega)$, there exist approximations $I_{h}(\textit{{w}})\in\mathbf{V}_{h}$ and $\rho_{h}(q)\in Q_{h}$, such that
    $|\textit{{w}}-I_{h}(\textit{{w}})|_{H^{1}(\Omega)}\leq Ch^{r}\|\textit{{w}}\|_{H^{2}(\Omega)},\quad\|q-\rho_{h}(q)\|_{L^{2}(\Omega)}\leq Ch^{r}\|q\|_{H^{1}(\Omega)},$
@@
There are several families of elements present in the literature that satisfy conditions A1 and A2 (see, for example, [Ar2018, AB2010, ABF1984, BDG2006, Boffi1995, BBDDFF2008] and the references therein). From now on, let us consider $\mathbf{g}_{\varepsilon}=\mathbf{g}_{h}$, i.e. a regular approximation of $\mathbf{g}$, such that $\mathbf{g}_{h}$ is the trace of a function in $\mathbf{V}_{h}$ and
@@
$\left\{\begin{array}[]{rll}a(\textit{{u}}_{h},\phi_{h})+c(\textit{{u}}_{h},\textit{{u}}_{h},\phi_{h})-b(\phi_{h},p_{h})&=\int_{\Omega}\mathbf{f}\cdot\phi_{h}\quad&\forall\phi_{h}\in\mathbf{V}_{h}\cap\mathbf{H}^{1}_{0}(\Omega)\\ b(\textit{{u}}_{h},q_{h})&=0\quad&\forall q_{h}\in Q_{h}\end{array}\right$. (17)
The existence and uniqueness of problem (17) is studied, for example, in [wang2016, gunzburger1983], for conforming finite element methods that satisfy A1 and A2, under appropriate conditions on the data.
In the next section, we consider some possible choices for the discrete spaces $\mathbf{V}_{h}$, $Q_{h}$ and for the function $\mathbf{g}_{h}$.
The following result provides an useful a priori error estimate for the finite element approximation of Navier-Stokes problem with regular data satisfying the conditions of Theorem 2.3.
@@
Let $\mathbf{f}\in\mathbf{H}^{-1}(\Omega)$. If A1 and A2 hold, we get the following error estimates
@@
Let $\mathbf{f}\in\mathbf{H}^{-1}(\Omega)$ be and $\mathbf{g}_{h}$ satisfying (1) and (16). Let $\textit{{u}}\in\mathbf{L}_{\Gamma}$ be the solution of the problem (2) and $\textit{{u}}_{h}$ be the solution of (17). If A1 and A2 hold, then, there exists a constant $C$ such that
@@ 5 Numerical experiments
In this section we discuss some numerical aspects of our approach and present numerical experiments for the classic lid-driven cavity flow problem, which is a typical benchmark problem in computational fluid dynamics [duran2020, hamouda2017, AB2010].
The first issue is to choose finite elements spaces $\mathbf{V}_{h}$ and $Q_{h}$ satisfying conditions A1 and A2. We will consider the following:
@@
  • $\mathbf{P}_{1}P_{1}$ stabilized: the method consists of find $(\textit{{u}}_{h},p_{h})\in\mathbf{V}_{h}\times Q_{h}$ with $\textit{{u}}_{h}=\mathbf{g}_{h}$ in $\Gamma$, $\mathbf{V}_{h}=(P_{1}(\mathcal{T}_{h}))^{2}$ and $Q_{h}=P_{1}(\mathcal{T}_{h})\cap L^{2}_{0}(\Omega)$ such that
    $\left\{\begin{array}[]{rll}a(\textit{{u}}_{h},\phi_{h})+c(\textit{{u}}_{h},\textit{{u}}_{h},\phi_{h})-b(\phi_{h},p_{h})&=\int_{\Omega}\mathbf{f}\cdot\phi_{h}&\forall\phi_{h}\in\mathbf{V}_{h}\cap\mathbf{H}^{1}_{0}(\Omega)\\ b(\textit{{u}}_{h},q_{h})-G(p_{h},q_{h})&=0&\forall q_{h}\in Q_{h}\end{array}\right$.
@@
    with $\Pi q|_{T}=\frac{1}{|T|}\int_{T}qdx,\forall T\in\mathcal{T}_{h}$. This method was proposed in [BDG2006] for the Stokes problem and studied for the Navier-Stokes problem in [zheng, HL2008].
After one of those methods is chosen, another question is how to select an appropriate function $\mathbf{g}_{h}$ that satisfies (16) and the compatibility condition (1). One possibility is to take $\mathbf{g}_{h}$ as the Cartensen interpolator of the function $\mathbf{g}$ [cartensen]. Another option is to take $\mathbf{g}_{h}$ as the Lagrange interpolator but, unfortunately, this choice doesn’t verify (1). In [duran2020, lemma 3.1] the authors propose a modification of the Lagrange interpolator so that (1) holds.
Let $\Omega=[-1,1]\times[-1,1]$, $\mathbf{f}=(0,0)$, $\nu=1$ and
@@
We will obtain the numerical approximations applying the three finite elements methods described above and choosing $\mathbf{g}_{h}$ as the Lagrange interpolator. We observe that, in this case, the compatibility condition is automatically verifed and furthermore, following [duran2020, Proposition 3.1], we have that
@@
Other numerical aspect to consider is the iterative methods to approximate the discrete solution. The typical algorithm to solve the Navier-Stokes problem is based on a fixed-point iterative algorithm. One possibility, that we will apply in our numerical experiments, is to use a Newton scheme, which has been studied for both the homogeneous [girault1986, karakashian1982] and non-homogeneous cases [wang2016, gunzburger1983].
@@
$a(\textit{{u}}^{n}_{h},\phi_{h})+c(\textit{{u}}^{n}_{h},\textit{{u}}^{n-1}_{h},\phi_{h})$ $+c(\textit{{u}}^{n-1}_{h},\textit{{u}}^{n}_{h},\phi_{h})-b(\phi_{h},p^{n}_{h})$
@@
In [gunzburger1983, Proposition 3.4 and Proposition 3.5] it is shown that if $\textit{{u}}^{0}_{h}$ is close enough to $\textit{{u}}_{h}$, then the sequence defined by (5) is unique and converges quadratically to $\textit{{u}}_{h}$. In other words, there exists a $\delta>0$ such that if $\|\textit{{u}}_{0}-\textit{{u}}_{h}\|_{H^{1}(\Omega)}\leq\delta$, then for $n\geq 1$,
@@
The numerical implementation is based on the $NGSolve$ Python package [NGSolve]. For the application of Newton’s method, a stopping criterion must be prescribed in terms of a given tolerance. In the present work, we employ the relative error between two consecutive iterates:
$\frac{|\textit{{u}}^{n}_{h}-\textit{{u}}^{n-1}_{h}|_{H^{1}(\Omega)}}{|\textit{{u}}^{n}_{h}|_{H^{1}(\Omega)}}<tol$
with $tol=10^{-8}$. As an initial guess for the Newton iteration, we take the solution of the corresponding Stokes problem.
@@
Figure 1 and table 1 show the graphics of $\log(h)$ vs $\log(e_{L^{4}})$ for $\mathbf{P}_{2}P_{1}$, $\mathbf{P}^{b}_{1}P_{1}$ and stabilized $\mathbf{P}_{1}P_{1}$ elements respectively, achieved with uniform mesh refinement, and also exhibit the experimental order of convergence (eoc) with respect to $h$, which is computed by comparing the errors for two consecutive meshes. We can observe that, in all the cases, the numerical results confirm the theorical prediction.
Figure 1: Convergence order analysis.
Table 1: $L^{4}$-norm errors and experimental orders of convergence for the approximation obtained with uniform mesh refinements, for different finite element discretizations.
$\mathbf{P}_{2}P_{1}$ $\mathbf{P}^{b}_{1}P_{1}$ $\mathbf{P}_{1}P_{1}$ stabilized
$h$ $e_{L^{4}}$ eoc $e_{L^{4}}$ eoc $e_{L^{4}}$ eoc
1/4 1.0116e-01 - 2.0160e-01 - 1.1894e-01 -
1/8 7.1461e-02 0.50146 1.4205e-01 0.50514 8.4146e-02 0.49931
1/16 5.0534e-02 0.49991 1.0044e-01 0.50006 5.9510e-02 0.49978
1/32 3.5729e-02 0.50015 7.1057e-02 0.49924 4.2082e-02 0.49992
1/64 2.5262e-02 0.50010 5.0247e-02 0.49995 2.9757e-02 0.49998
[…]